OpenStudy (anonymous):

A 6 ft adult has a shadow 3.6 ft long. How long is the shadow of a 5 ft child standing next to the adult?

OpenStudy (anonymous):



OpenStudy (anonymous):

Okay, this problem involves cross multiplying. 6 5 --- = --- 3.6 x 6 * x = 6x 3.6 * 5 = 18 Now, you divide. 6x = 18 6 cancels out. What you do on one side, you do to the other, so now it's: 18 / 6 = 3 x = 3 So, if a 6 foot tall adult has a shadow of 3.6 feet long, then a child that is 5 feet tall has a shadow of 3 feet.

OpenStudy (anonymous):

\[\Large \begin{array}{l} \tan \theta = \frac{o}{a}\\ {\tan ^{ - 1}}\frac{{3.6}}{6} = 30.96^\circ \\ o = \tan \theta \cdot a = \tan 30.96 \cdot 5 \approx 3 \end{array}\]